# Yufei Bo, Shuo Shao, Meixia Tao

### v1

@@ V Conclusion
This paper proposes a new framework for digital semantic communications over degraded AWGN broadcast channels, namely the DeepSCM scheme. In this scheme, the semantic features intended for different receivers are encoded into a basic encoded feature vector and its successive refinement vector, which are then associated with different layers of a super-constellation. To minimize redundancy in broadcasting, an LMMSE decorrelator is developed to ensure that these two vectors are nearly uncorrelated with each other. This superposition code structure can accommodate the communication requirements of different receivers with diverse channel conditions. The proposed scheme is especially effective in scenarios with large channel disparity and high modulation order. All in all, the proposed framework not only provides an efficient way to conduct semantic broadcasting, but also shows a promising future of combining theoretical coding schemes with NN-based coding method.
In the future, we plan to extend our superposition coding approach to other multi-user communication scenarios, such as the semantic communications for multi-terminal sources. Moreover, we will enhance the performance of our proposed scheme by introducing techniques such as superposition constellation shaping.
@@ Appendix A Proof of Proposition 1
[…]
$\overset{(d)}{\leq}n\log\frac{\pi e}{n}\sum_{i=1}^{2n}\mathbf{E}[R_{i}^{2}]$
$=n\log\frac{\pi e}{n}\mathbf{E}[||\mathbf{R}||^{2}_{2}],$ (21)
[…]

### v2

@@ V Conclusion
This paper proposes a new framework for digital semantic communications over degraded AWGN broadcast channels, namely the DeepSCM scheme. In this scheme, the semantic features intended for different receivers are encoded into a basic EFV and its successive refinement vector, which are then associated with different layers of a super-constellation. To minimize redundancy in broadcasting, an LMMSE decorrelator is developed to ensure that these two vectors are nearly uncorrelated with each other. This superposition code structure can accommodate the communication requirements of different receivers with diverse channel conditions. The proposed scheme is especially effective in scenarios with large channel disparity and high modulation order. This work not only provides an efficient way to conduct semantic broadcasting, but also showcases a promising approach of combining theoretical coding schemes with NN-based coding method.
@@ Appendix A Proof of Proposition 1
$h(\mathbf{R})$ $\overset{(a)}{\leq}\frac{1}{2}\log\left\{(2\pi e)^{2n}|\textit{Var}[\mathbf{R}]|\right\}$
$\overset{(b)}{\leq}\frac{1}{2}\log\left\{(2\pi e)^{2n}\textstyle\prod_{i=1}^{2n}\textit{Var}[R_{i}]\right\}$
$\overset{(c)}{\leq}\frac{1}{2}\log\left\{(2\pi e)^{2n}(\frac{\sum_{i=1}^{2n}\textit{Var}[R_{i}]}{2n})^{2n}\right\}$
$\overset{(d)}{\leq}n\log\left\{\frac{\pi e}{n}\sum_{i=1}^{2n}\mathbf{E}[R_{i}^{2}]\right\}$
$=n\log\left\{\frac{\pi e}{n}\mathbf{E}[||\mathbf{R}||^{2}_{2}]\right\},$ (23)
where $R_{i}$ denotes the $i$th entry of $\mathbf{R}$, $|\textit{Var}[\mathbf{R}]|$ denotes the determinant of $\textit{Var}[\mathbf{R}]$, $(a)$ follows from the fact that Gaussian distribution maximizes entropy under a given covariance matrix, $(b)$ follows from Hadamard inequality since covariance matrices are always positive semidefinite, $(c)$ follows from the inequality of arithmetic and geometric means, and $(d)$ follows from the definition of variance that $\textit{Var}[R_{i}]=\mathbf{E}[R_{i}^{2}]-\mathbf{E}^{2}[R_{i}]$.
We have equality in $(a)$ if $\mathbf{R}$ is Gaussian distributed; we have equality in $(b)$ if the elements of $\mathbf{R}$ are mutually uncorrelated; the equality in $(c)$ holds if $\textit{Var}[R_{1}]=\textit{Var}[R_{2}]=...=\textit{Var}[R_{2n}]$ according to the inequality of arithmetic and geometric means; and the equality in $(d)$ holds if all elements in $\mathbf{R}$ have zero mean. Combining these facts together suggests the equality holds if each element of $\mathbf{R}$ follows an i.i.d. Gaussian distribution with zero mean.